# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v2

@@ Abstract
In this work, we propose Wasserstein distributionally robust shallow convex neural networks (WaDiRo-SCNNs) to provide reliable nonlinear predictions when subject to adverse and corrupted datasets. Our approach is based on a new convex training program for $\operatorname{ReLU}$-based shallow neural networks which allows us to cast the problem as an exact, tractable reformulation of its order-1 Wasserstein distributionally robust counterpart. Our training procedure is conservative, has low stochasticity, is solvable with open-source solvers, and is scalable to large industrial deployments. We provide out-of-sample performance guarantees, show that hard convex physical constraints can be enforced in the training program, and propose a mixed-integer convex post-training verification program to evaluate model stability. WaDiRo-SCNN aims to make neural networks safer for critical applications, such as in the energy sector. Finally, we numerically demonstrate the performance of our model on a synthetic experiment, a real-world power system application, i.e., the prediction of non-residential buildings’ hourly energy consumption in the context of virtual power plants, and on benchmark datasets. The experimental results are convincing and showcase the strengths of the proposed model.
Keywords: distributionally robust optimization, shallow convex neural networks, trustworthy machine learning, physics constrained, energy.
@@ 1 Introduction
Consider the following general feedforward neural network (NN) training problem for a network of $h\in\mathbb{N}$ hidden layers under a loss function $\mathcal{L}:\mathbb{R}^{N\times m_{h+1}}\times\mathbb{R}^{N\times m_{h+1}}\to\mathbb{R}$ which penalizes the difference between the $N\in\mathbb{N}$ training labels and NN predictions:
@@
where $\phi_{i}:\mathbb{R}^{m_{i}}\to\mathbb{R}^{m_{i+1}}$ is the activation function of the hidden layer $i$, $m_{i}\in\mathbb{N}^{*}$ is the number of neurons in layer $i$, $\mathbf{W}_{i}\in\mathbb{R}^{m_{i-1}\times m_{i}}$ are the weights in layer $i$, and $\mathbf{b}_{i}\in\mathbb{R}^{m_{i}}$ are the bias weights of layer $i$ for all $i\in\{1,2,3,\ldots,h\}=\llbracket h\rrbracket$. Let $\mathbf{x}_{j}\in\mathbb{R}^{m_{0}},\mathbf{y}_{j}\in\mathbb{R}^{m_{h+1}},\ \text{and}\ \hat{\mathbf{y}}_{j}\in\mathbb{R}^{m_{h+1}}$ define training features, training labels, and training predictions, respectively, $\forall j\in\llbracket N\rrbracket$. Matrices $\mathbf{Y},\hat{\mathbf{Y}}\in\mathbb{R}^{N\times m_{h+1}}$ are the collection of all labels and predictions. In our notation, we assume that NN inputs and outputs are vectors. A common activation function $\phi_{i}$ for regression is the element-wise rectified linear unit ($\operatorname{ReLU}$) operator:
$\operatorname{ReLU}(\cdot)=\max\{\mathbf{0},\cdot\},$
@@
as it is accurate when modelling nonlinear patterns and biomimics the required excitation threshold for brain neurons to activate. The acquisition of features is done through the input layer and the output of labels through the output layer such that a neural network of $h$ hidden layers has effectively $h+2$ layers. To solve (NNT), the loss is propagated iteratively, one layer at a time, from the output neurons to the input neurons, by computing gradient estimators of the NN architecture at each layer and updating weights with methods derived from non-convex optimization, e.g., stochastic gradient descent (Amari 1993). This procedure is referred to as backpropagation (Bishop and Bishop 2023).
@@
In part because (NNT) is typically non-convex and only solvable through stochastic heuristics with many hyperparameters, NNs have been traditionally disregarded in critical sectors, e.g., energy (Chatzivasileiadis et al. 2022), healthcare (Rasheed et al. 2022), and finance (Giudici and Raffinetti 2023). Even though they tend to be successful nonlinear predictors, their lack of interpretability, limited out-of-the-box performance guarantees, significant data requirements, as well as high susceptibility to data corruption and other adversarial attacks, have labelled them as unreliable for such applications.
@@
In the past decade, many solutions have been proposed to guide and certify NN training as much as possible to mitigate these inherent flaws before deployment. For example, blackbox optimization (BBO) algorithms with provable convergence properties can add some local convergence guarantees in the hyperparameter tuning phase (Audet et al. 2022, Kawaguchi and Sun 2021) because NNs’ loss landscape, with respect to their hyperparameters, can be fractal (Sohl-Dickstein 2024). However, when paired with highly stochastic training programs, their relevance is diminished.
@@
Because their high complexity and opacity make them unpredictable, complete performance assessment is hard to obtain during training. As such, post-training verification frameworks have emerged as a way to certify formal guarantees on trained machine learning (ML) models (Huang et al. 2017). These procedures are conceptually similar to quality control tests in factories. By using a model’s trained weights as parameters and representing the model’s internal mechanisms in an optimization problem, these frameworks aim to find theoretical worst-case scenarios, e.g., worst-case constraint violation in a specific critical application (Venzke et al. 2020), worst-case instability (Huang et al. 2021), and more (Albarghouthi 2021). A general feedforward neural network verification problem takes the form:
$\max_{\mathbf{x},\hat{\mathbf{y}},\bm{\sigma}}$ $\mathcal{V}\left(\hat{\mathbf{y}},\mathbf{x},\bm{\sigma}\right)$ (PTV)
@@
where $\mathbf{x}\in\mathcal{X}$ is any possible input from the feature space $\mathcal{X}\subseteq\mathbb{R}^{m_{0}}$, $\mathcal{V}:\mathbb{R}\times\mathbb{R}^{m_{h+1}}\times\mathbb{R}^{\operatorname{card}(\bm{\sigma})}\to\mathbb{R}$ is an objective function which measures a specific guarantee, $\bm{\sigma}$ represents all of the problem’s other variables used in the verification process, and $f_{1},f_{2},\dots,f_{n},g_{1},g_{2},\dots,g_{l}$ are the family of inequality and equality constraints needed for the certification. Note that a convex representation of the internal mechanisms with respect to the inputs is extremely desirable for such problems because the convergence guarantees to global optimality of convex optimization eliminate any doubt on the verification’s validity.
@@
When a neural network is used in a dynamical setting, i.e., a system that evolves with time over a state-space and follows a set of physical rules (Wiggins 2003), the introduction of physical knowledge from the setting into its training procedure can increase its overall reliability and greatly reduce the data requirements to obtain good empirical performances (Karniadakis et al. 2021, Xu et al. 2023). This type of training is often referred to as physics-informed, physics-guided, or physics-constrained depending on how physical knowledge is embedded in the training process. We refer interested readers to Cuomo et al. (2022) for an overview of different techniques. We will cover the most common one in Section 1.2.
@@
Typically, as it is hard to guarantee NN out-of-sample performance over time and their performance may degrade (drift), the aforementioned techniques are combined in complex pre-deployment and post-deployment loops (Stiasny et al. 2022) with tight MLOps (contraction of machine learning, software development, and operations) procedures for lifecycle management (Kreuzberger et al. 2023). Good MLOps practices help monitor and assert, theoretically and empirically, the quality of ML models from training to deployment but require great computational infrastructure to be well implemented. Such infrastructures are not democratized in every sector, especially energy and healthcare. We present an example of a safe MLOps pipeline inspired by Stiasny et al. (2022) in Figure 1. A safe MLOps pipeline can be divided into three parts: (i) definition of initial conditions, (ii) hyperparameter optimization, and (iii) model verification. Initial conditions englobe the choice of dataset, ML model, hyperparameter tuning algorithm, and the definition of the hyperparameter search space. Hyperparameter optimization covers the typical training and validation phases: hyperparameters are chosen for each run to pre-process the training dataset and train the ML model, the trained model is then evaluated on the validation dataset, and different metrics are logged. This part ends when a stopping criterion is met, e.g., the maximum number of runs is reached. Finally, the best models in the validation phase are verified through post-training certification methods and evaluated on the testing dataset. If the performance expectations are not met by any model, initial conditions must be redefined. Otherwise, the best ML model can be deployed. Most of the computational burden stems from hyperparameter optimization. As such, there is a growing interest in designing complex predictive models with low-stochasticity training procedures, e.g., few hyperparameters and convex with respect to the weights, as well as lower data pre-processing requirements, e.g., by being distributionally robust. These models simplify the hyperparameter optimization phase and have more chance than standard models to meet the requirements of the verification phase.
@@
In this work, our motivation is to design a nonlinear predictive ML model with a simple low-stochasticity training program that can be guided by physical constraints and that has inherent performance guarantees. We also aim to design this model with an architecture that can be easily included in post-training verification frameworks or data-driven control applications, if needs be. Shallow convex neural networks, as we will show, are perfect candidates for the task. In doing so, we hope to lower computational infrastructure needs for safe deployment in large-scale critical sectors where a lack of expertise and budget constraints do not permit state-of-the-art MLOps pipelines at all stages, and yet accurate nonlinear predictions are required. We specifically target virtual power plants (VPPs) (Pudjianto et al. 2007) applications where reliable forecasts of distributed assets are crucial for the safe operations of the electrical power grid.
As an example of a standard VPP application, we utilize our model to predict the hourly energy consumption of non-residential buildings enrolled in demand response programs (Siano 2014). Because of their immense impact on distribution grids and their diverse activities, non-residential buildings have complex nonlinear consumption patterns that must be predicted accurately to enhance grid operation (Massana et al. 2015). As each building needs its model and thus possibly thousands of models must be deployed, scalability issues arise when utilizing complex stochastic ML training procedures in complete safe MLOps pipelines. Finally, the sensitivity of the application, and the risk associated with poor predictions, promote models with theoretical guarantees, mechanisms to respect physical constraints, and robustness to possible anomalies in the training data, e.g., numerical errors, noise in sensor readings, telemetry issues, meters outage, and unusual extreme events (Zhang et al. 2021).
@@ 1.1 Notation
  • Bold lowercase letters refer to vectors, bold uppercase letters refer to matrices, and normal lowercase or uppercase letters refer to scalars. Bold Greek letters may refer to matrices depending on context.
  • The sets of real, natural, and integer numbers are denoted by $\mathbb{R}$, $\mathbb{N}$, and $\mathbb{Z}$, respectively.
  • Calligraphic uppercase letters refer to sets, e.g., $\mathcal{X}$ and $\mathcal{Z}$, or functions depending on context, e.g., $\mathcal{L}(\cdot)$ and $\mathcal{V}(\cdot)$.
@@
  • Let $\mathbf{a}\in\mathbb{R}^{n}$ be a general vector of dimension $n\in\mathbb{N}^{*}$, column vectors are compactly written $\mathbf{a}=\left(a_{1},a_{2},\ldots,a_{n}\right)$ and row vectors as $\mathbf{a}^{\top}=\left(a_{1},a_{2},\ldots,a_{n}\right)^{\top}=\left(a_{1}\ a_{2}\ \ldots\ a_{n}\right)$.
[…]
@@
  • $\operatorname{vec}(\cdot):\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$ denotes the vectorization operator which flattens matrices. If $\mathbf{A}=(\mathbf{a}_{1}\ \mathbf{a_{2}}\ \ldots\ \mathbf{a_{n}})$ where $\mathbf{a}_{i}\in\mathbb{R}^{m},\ n,m\in\mathbb{N}^{*},\ \forall i\in\llbracket n\rrbracket$, then $\operatorname{vec}(\mathbf{A})=(\mathbf{a}_{1}^{\top}\ \mathbf{a}_{2}^{\top}\ \ldots\ \mathbf{a}_{n}^{\top})^{\top}$.
[…]
  • $\mathbf{0}_{n}\in\mathbb{R}^{n}$ refers to the null vector of size $n$.
  • $\ell_{1}$-norm refers to $\|\cdot\|_{1}$, $\ell_{2}$-norm refers to $\|\cdot\|_{2}$, $\ell_{1}$-loss refers to $\|\mathbf{Y}-\hat{\mathbf{Y}}\|_{1}$, and $\ell_{2}$-loss refers to $\|\mathbf{Y}-\hat{\mathbf{Y}}\|_{2}^{2}$.

### v3

@@ Abstract
In this work, we propose Wasserstein distributionally robust shallow convex neural networks (WaDiRo-SCNNs) to provide reliable nonlinear predictions when subject to adverse and corrupted datasets. Our approach is based on the reformulation of a new convex training program for $\operatorname{ReLU}$-based shallow neural networks, which allows us to cast the problem into the order-1 Wasserstein distributionally robust optimization framework. Our training procedure is conservative, has low stochasticity, is solvable with open-source solvers, and is scalable to large industrial deployments. We provide out-of-sample performance guarantees, show that hard convex physical constraints can be enforced in the training program, and propose a mixed-integer convex post-training verification program to evaluate model stability. WaDiRo-SCNN aims to make neural networks safer for critical applications, such as in the energy sector. Finally, we numerically demonstrate our model’s performance through both a synthetic experiment and a real-world power system application, viz., the prediction of hourly energy consumption in non-residential buildings within the context of virtual power plants, and evaluate its stability across standard regression benchmark datasets. The experimental results are convincing and showcase the strengths of the proposed model.
Keywords: Distributionally robust optimization, Shallow convex neural networks, Trustworthy machine learning.
@@ 1 Introduction
Consider a general feedforward neural network (NN) training problem for a network of $h\in\mathbb{N}$ hidden layers under a loss function $\mathcal{L}:\mathbb{R}^{N\times m_{h+1}}\times\mathbb{R}^{N\times m_{h+1}}\to\mathbb{R}$ which penalizes the difference between the $N\in\mathbb{N}$ training labels and NN predictions of dimension $m_{h+1}$, the number of hidden neurons in the output layer. Let $\mathbf{x}_{j}\in\mathbb{R}^{m_{0}},\mathbf{y}_{j}\in\mathbb{R}^{m_{h+1}},\ \text{and}\ \hat{\mathbf{y}}_{j}\in\mathbb{R}^{m_{h+1}}$ define training features, training labels, and training predictions, respectively, $\forall j\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{1,2,3,\ldots,N\}\eqqcolon}}\llbracket N\rrbracket$. Matrices $\mathbf{Y},\hat{\mathbf{Y}}\in\mathbb{R}^{N\times m_{h+1}}$ denote the collection of all labels and predictions. The problem is as follows:
@@
where ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi^{i}}}:\mathbb{R}^{m_{i}}\to\mathbb{R}^{m_{i+1}}$ is the activation function of the hidden layer $i$, $m_{i}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{N}_{>0}}}$ is the number of neurons in layer $i$, $\mathbf{W}_{i}\in\mathbb{R}^{m_{i-1}\times m_{i}}$ are the weights in layer $i$, and $\mathbf{b}_{i}\in\mathbb{R}^{m_{i}}$ are the bias weights of layer $i$ for all $i\in\llbracket h\rrbracket$. In our notation, we assume that NN inputs and outputs are vectors. Let $\mathbf{0}_{m_{i}}$ be the null vector of dimension $m_{i}$, a common activation function $\phi^{i}$ for regression is the element-wise rectified linear unit ($\operatorname{ReLU}$) operator:
$\operatorname{ReLU}(\cdot)=\max\{\mathbf{0}_{m_{i}},\cdot\},$
@@
as it is accurate when modelling nonlinear patterns and biomimics the required excitation threshold for brain neurons to activate (Householder 1941). The acquisition of features is done through the input layer while the resulting labels are obtained through the output layer. Thus, a neural network of $h$ hidden layers has effectively $h+2$ layers. To solve (NNT), the gradient of the loss is propagated iteratively, one layer at a time, from the output neurons to the input neurons, by computing gradient estimators of the NN architecture at each layer and updating weights with methods derived from non-convex optimization, e.g., stochastic gradient descent (SGD) (Amari 1993). This procedure is referred to as backpropagation (Bishop and Bishop 2023).
@@
In part because (NNT) is typically non-convex and only solvable through stochastic heuristics with many hyperparameters, NNs have traditionally been disregarded in critical sectors, e.g., energy (Chatzivasileiadis et al. 2022), healthcare (Rasheed et al. 2022), and finance (Giudici and Raffinetti 2023). Even though they tend to be successful nonlinear predictors, their lack of interpretability, limited out-of-the-box performance guarantees, significant data requirements, as well as high susceptibility to data corruption and other adversarial attacks, have labelled them as unreliable for such applications.
In the past decade, many solutions have been proposed to guide and certify NN training as much as possible to mitigate these inherent flaws before deployment.
@@
As a first example, blackbox optimization (BBO) algorithms with provable convergence properties can provide some local convergence guarantees in the hyperparameter tuning phase (Audet et al. 2022, Kawaguchi and Sun 2021) because NNs’ loss landscape, with respect to their hyperparameters, can be fractal (Sohl-Dickstein 2024). However, when paired with highly stochastic training programs, their relevance is diminished.
@@
Because their high complexity and opacity make them unpredictable, complete performance assessment is hard to obtain during training. As such, post-training verification frameworks have emerged as a way to certify formal guarantees on trained machine learning (ML) models (Huang et al. 2017). These procedures are conceptually similar to quality control tests in factories. By using a model’s trained weights as parameters and representing the model’s internal mechanisms in an optimization problem, these frameworks aim to find theoretical worst-case scenarios, e.g., worst-case constraint violation in a specific critical application (Venzke et al. 2020), worst-case instability (Huang et al. 2021), and more (Albarghouthi 2021). A general feedforward neural network post-training verification (PTV) problem takes the form:
$\max_{\mathbf{x},\hat{\mathbf{y}},\bm{\sigma}}\$$\mathcal{V}\left(\hat{\mathbf{y}},\mathbf{x},\bm{\sigma}\right)$ (PTV)
@@
where $\mathbf{x}\in\mathcal{X}$ is any possible input from the (bounded) feature space $\mathcal{X}\subseteq\mathbb{R}^{m_{0}}$, $\mathcal{V}:\mathbb{R}\times\mathbb{R}^{m_{h+1}}\times\mathbb{R}^{\operatorname{card}(\bm{\sigma})}\to\mathbb{R}$ is an objective function which measures a specific guarantee, $\bm{\sigma}$ represents all of the problem’s other variables used in the verification process, and $f_{1},f_{2},\dots,f_{n},g_{1},g_{2},\dots,g_{l}$ are the family of inequality and equality constraints needed for the certification. Note that a convex representation of the internal NN mechanisms, with respect to the inputs, is desirable for such problems because the convergence guarantees to global optimality of convex optimization eliminate any doubt regarding the validity of the verification.
@@
When a neural network is used in a dynamical setting, in which a system evolves over time on a state-space and follows a set of physical rules (Wiggins 2003), the introduction of physical knowledge from the setting into its training procedure can increase its overall reliability and greatly reduce the data requirements to obtain good empirical performances (Karniadakis et al. 2021, Xu et al. 2023). This type of training is often referred to as physics-informed, physics-guided, or physics-constrained depending on how physical knowledge is embedded in the training process. We refer interested readers to Cuomo et al. (2022) for an overview of different techniques. We will cover the most common one in Section 1.2.
@@
Typically, as it is hard to guarantee NN out-of-sample performance over time and their performance may degrade (drift), the aforementioned techniques are combined in complex pre-deployment and post-deployment loops (Stiasny et al. 2022) with tight MLOps (contraction of machine learning, software development, and operations) procedures for lifecycle management (Kreuzberger et al. 2023). Extensive MLOps practices help monitor and assert, theoretically and empirically, the quality of ML models from training to deployment, but require substantial computational infrastructures to be well implemented. Such infrastructures are not democratized in every sector, especially energy and healthcare. We present an example of a safe MLOps pipeline inspired by Stiasny et al. (2022) in Figure 1. A safe MLOps pipeline can be divided into three parts: (i) definition of initial conditions, (ii) hyperparameter optimization, and (iii) model verification. Initial conditions englobe the choice of dataset, ML model, hyperparameter tuning algorithm, and the definition of the hyperparameter search space. Hyperparameter optimization covers the typical training and validation phases: hyperparameters are chosen for each run to pre-process the training dataset and train the ML model, the trained model is then evaluated on the validation dataset, and different metrics are logged. This part ends when a stopping criterion is met, e.g., the maximum number of runs is reached. Finally, the best models in the validation phase are verified through post-training certification methods and evaluated on the testing dataset. If the performance expectations are not met by any model, initial conditions must be redefined. Otherwise, the best ML model can be deployed.
Most of the computational burden stems from hyperparameter optimization. As such, there is a growing interest in designing complex predictive models with low-stochasticity training procedures, e.g., involving fewer hyperparameters and exhibiting convexity with respect to the weights. There is also an interest in designing training procedures that can withstand imperfect training data and thus have lower data pre-processing requirements, e.g., by being distributionally robust. These models simplify the hyperparameter optimization phase and are more likely than standard models to meet the requirements of the verification phase.
@@
In this work, our motivation is to design a nonlinear predictive ML model with a simple low-stochasticity training program that can be guided by physical constraints and that has inherent performance guarantees. We also aim to design this model with an architecture that can be easily included in post-training verification frameworks or data-driven control applications, if required. Shallow convex neural networks, as we will show, are well-suited for the task. In doing so, we hope to lower computational infrastructure needs for safe deployment in large-scale critical sectors where a lack of expertise and budget constraints do not permit state-of-the-art MLOps pipelines at all stages, and yet accurate nonlinear predictions are required. We specifically target virtual power plants (VPPs) (Pudjianto et al. 2007) applications where reliable forecasts of distributed assets are crucial for the safe operations of the electrical power grid.
As an example of a standard VPP application, we utilize our model to predict the hourly energy consumption of non-residential buildings enrolled in demand response programs (Siano 2014). Because of their significant impact on distribution grids and their diverse activities, non-residential buildings have complex nonlinear consumption patterns that must be predicted accurately to enhance grid operation (Massana et al. 2015). As each building needs its model and thus possibly thousands of models must be deployed, scalability issues arise when utilizing complex stochastic ML training procedures in complete safe MLOps pipelines. Finally, the sensitivity of the application, and the risk associated with poor predictions, promote models with theoretical guarantees, mechanisms to respect physical constraints, and robustness to possible anomalies in the training data, e.g., numerical errors, noise in sensor readings, telemetry issues, meters outage, and unusual extreme events (Zhang et al. 2021).
@@ 1.1 Notation
  • Bold lowercase letters, e.g., $\mathbf{a}$, refer to vectors, bold uppercase letters, e.g., $\mathbf{A}$, refer to matrices, and normal lowercase or uppercase letters refer to scalars, e.g., $a$, or $A$. Bold Greek letters may refer to matrices depending on context.
  • The sets of real, natural, and integer numbers are denoted by $\mathbb{R}$, $\mathbb{N}$, and $\mathbb{Z}$, respectively. We add the subscript $\{\cdot\}_{>0}$, e.g., $\mathbb{N}_{>0}$, to represent the respective subset of strictly positive numbers.
  • We use the following combinations of font style and letters for distributions: $\text{D},\text{F},\text{P},\text{Q},\text{U}$, and V.
  • Calligraphic uppercase letters refer to sets, e.g., $\mathcal{X}$ and $\mathcal{Z}$, or to functions depending on context, e.g., $\mathcal{L}(\cdot)$ and $\mathcal{V}(\cdot)$.
  • Letters $\mathbf{x},\ \mathbf{y},\text{ and }\mathbf{z}$ are reserved for features, labels, and samples, respectively, while $\hat{\mathbf{y}}$ refers to model predictions.
  • Let $\mathbf{a}\in\mathbb{R}^{n}$ be a general vector of dimension $n\in\mathbb{N}_{>0}$, column vectors are compactly written $\mathbf{a}=\left(a_{1},a_{2},\ldots,a_{n}\right)$ and row vectors as $\mathbf{a}^{\top}=\left(a_{1},a_{2},\ldots,a_{n}\right)^{\top}=\left(a_{1}\ a_{2}\ \ldots\ a_{n}\right)$.
@@
  • $\{\cdot\}\succeq 0$ refers to semidefinite positiveness if $\{\cdot\}$ is a matrix and is an element-wise $\geq 0$ if $\{\cdot\}$ is a vector.
@@
  • $\operatorname{vec}(\cdot):\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$ denotes the vectorization operator which flattens matrices. If $\mathbf{A}=(\mathbf{a}_{1}\ \mathbf{a_{2}}\ \ldots\ \mathbf{a_{n}})$ where $\mathbf{a}_{i}\in\mathbb{R}^{m},\ n,m\in\mathbb{N}_{>0},\ \forall i\in\llbracket n\rrbracket$, then $\operatorname{vec}(\mathbf{A})=(\mathbf{a}_{1}^{\top}\ \mathbf{a}_{2}^{\top}\ \ldots\ \mathbf{a}_{n}^{\top})^{\top}$.
  • $\operatorname{card}(\{\cdot\})$ refers to the cardinality of the set $\{\cdot\}$.
  • $\mathbf{0}_{n}\in\mathbb{R}^{n}$ refers to the null vector of size $n\in\mathbb{N}_{>0}$.
  • $\|\cdot\|_{1}$ refers to the $\ell_{1}$-norm , $\|\cdot\|_{2}$ refers to the $\ell_{2}$-norm, $\|\mathbf{y}-\hat{\mathbf{y}}\|_{1}$ refers to the $\ell_{1}$-loss, and $\|\mathbf{y}-\hat{\mathbf{y}}\|_{2}^{2}$ refers to the $\ell_{2}$-loss where $\mathbf{y}$ and $\hat{\mathbf{y}}$ are the vectors of all training labels and model predictions, respectively.